OpenStudy (anonymous):

Determine the number of real solutions of -2x^2 + 5x - 3 = 0. A. 0 B. 1 (double root) C. 2

OpenStudy (anonymous):

NOt sure how to do this

Directrix (directrix):

What is the discriminant?

OpenStudy (anonymous):

ax^+bx+c

Directrix (directrix):

Attached are two charts to help you think about this type problem.

OpenStudy (anonymous):

ax^2*

Directrix (directrix):

ax^2+bx+c is not the discriminant. Read about the discriminant on the attachment above.

OpenStudy (anonymous):

b^2-4ac

OpenStudy (anonymous):

ok let me try that

OpenStudy (anonymous):

is it B

Directrix (directrix):

I got 1 for the discriminant but that does not mean that there is a double root.

OpenStudy (anonymous):

Ok.

Directrix (directrix):

Look at the chart above with the title discriminant and root types. Under the discriminant column, find the description that fits the type number 1 is and read across the chart and look at the root description. Do that now, please. @Reaper534

OpenStudy (anonymous):

2 irrational roots?

OpenStudy (anonymous):

Wait 1 is a perfect square right so it would be 2 rational roots right?

OpenStudy (anonymous):

\[-2x^2 + 5x - 3 = 0.\]easier to solve \[2x^2-5x+3=0\] use the quadratic formula

OpenStudy (anonymous):

@satellite73 i got -6 and 6.5

OpenStudy (anonymous):

-6.5*

OpenStudy (anonymous):

i don't think that is right

OpenStudy (anonymous):

this one factors as \[2x^2-5x+3=(2x-3)(x-1)\]

OpenStudy (anonymous):

@satellite73 would it be equal to zero?

OpenStudy (anonymous):

like (2x-3)(x+1)=0

Directrix (directrix):

@Reaper534 >>>Wait 1 is a perfect square right so it would be 2 rational roots right? Correct

OpenStudy (anonymous):

Thanks! So its C correct?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

C) 2 is correct @Reaper534